# TaoSearchEmb: A Multi-Objective Reinforcement Learning Framework for Dense Retrieval in Taobao Search

### v1

@@ 4.1.1. Dataset and Metrics
For offline evaluation, we construct three test sets: the general retrieval test set, long-tail query evaluation set, and long-tail experience evaluation set, all collected and filtered from Taobao’s online search logs. The model is evaluated using Hitrate@6k and Goodrate@100.
@@
Next, we provide a concise description of the three offline test sets:
@@ 4.1.2. Implementation Details
The model is trained on a cluster of 256 GPUs (each equipped with 96GB memory). The training framework is implemented based on ROLL, an open-source distributed RL-training library optimized for large language models. During the Supervised Fine-Tuning (SFT) phase, a per-GPU batch size of 256 is configured with the AdamW optimizer ($\beta_{1}$=0.9, $\beta_{2}$=0.95), a learning rate of 1e-5, and a cosine learning rate scheduler incorporating a warmup ratio of 0.05. The architecture enforces a maximum input length of 256 tokens, with query and item embeddings projected into a 1024-dimensional L2-normalized space. MRL is applied across dimensions [1024, 2048, 3072] to enhance hierarchical representation learning. This phase completes after 1 full training epoch over the dataset.
@@ 4.1.3. Baselines
  • Query-Rewrite: Query rewriting via CSA-QR(feng2025complicated), followed by retrieval via inverted index.
  • Qwen-3B(yang2025qwen3): Dense retrieval model trained based on Qwen-3B, modifying to bi-directional attention.
  • Tbstars Post-Pretrain(base): Dense retrieval model trained based on an internal close-source Tbstars with 3B parameters, modifying to bi-directional attention.
@@
Table 3. Online Side-by-Side Human Evaluations
Query Type Case GSB Query Goodrate Item Goodrate
Q&A What medicine can make hair black? +24.30% +10.64pt +8.2pt
Alternative Alternative PS5 controller +28.57% +10.87pt +9.1pt
Overall - +29.88% +12.81pt +9.14pt
@@ 4.2. Offline Evaluation
As shown in Table 1, the model with SFT and Retrieval-GRPO achieves optimal performance, outperforming the state-of-the-art baseline Tbstars (SFT only) with an average improvement of 4 pt across four challenging query slices. The most significant enhancement is observed in the alternative querys, showing a 4.62 pt gain. It because the alternatives typically share identical functionality with original items but differ in brand and price. From the lexical overlap metric perspective, queries exhibit higher overlap scores with original brands than their alternatives, bringing challenges for models relying primarily on lexical matching. For instance, a search for ”alternative PS5 controller” might incorrectly prioritize official Sony products due to brand name overlap, despite the user’s explicit request for alternatives. Through Retrieval-GRPO, wrong predictions can be corrected via relevance rewards, which helps the model overcome superficial lexical matches and develop deeper semantic understanding.
@@
4.3. Ablation Study
4.3.1. Effect of R-GRPO
As shown in Table 2, we conduct ablation studies on the two-stage training framework. The SFT+Retrieval-GRPO approach significantly outperforms the SFT-only baseline. Additionally, replacing the second stage with ranking paradigm or DPO training with offline hard negative samples both resulted in performance degradation, demonstrating the clear advantages of our proposed Retrieval-GRPO phase. Furthermore, when we removed the first-stage SFT and directly applied reinforcement learning, the performance declined substantially, indicating that while Retrieval-GRPO is effective, the first SFT stage remains essential. The training framework can be viewed as a progressive difficulty training approach: if the model lacks the basic ability to distinguish between simple positive and negative samples, it will struggle even more with challenging samples.
4.3.2. Effect of Multi-Objective Reward
4.3.3. Effect of Reward Model
@@ 4.4. Online Evaluation
We deploy the model into the real-world pipeline and validated its effectiveness in the online search engine through A/B testing. A side-by-side evaluation was conducted between two experimental buckets, where the top 10 items for 2000 actual online query across four challenging query slices are manually annotated using three human evaluation metrics introduced in subsection 4.1.
The specific results are shown in Table 3. We conducted fine-grained analysis on four selected challenging query slices. The most significant improvement was observed in negative-class queries, as negative semantics and literal overlap metrics typically exhibit opposing trends, posing additional challenges for models that rely on literal matching rather than deep semantic understanding for traditional dense retrieval model. Through Retrieval-GRPO, the model leverages the semantic generalization capabilities of relevance models based on LLM, and employs real-time relevance feedback on retrieved top-items via reinforcement learning during training, enabling timely correction of wrong predictions.
@@ 5. Conclusion
This paper presents Retrieval-GRPO, a multi-objective reinforcement learning framework designed to address the limitations of conventional dense retrieval paradigms in e-commerce search systems. By leveraging dynamic candidate retrieval and multi-objective reward fusion in RL paradigm, our framework eliminates the dependency on labor-intensive hard negative sample mining while mitigating the seesaw effect inherent in multi-objective optimization. The proposed method integrates real-time relevance feedback from an LLM-based reward model, item quality metrics and exclusivity constraints, and dynamically optimizes semantic representations through reinforcement learning. Extensive experiments demonstrate that Retrieval-GRPO significantly enhances semantic generalization capabilities, particularly for complex long-tail queries, by enabling real-time error correction and preference alignment during training. The framework’s elimination of static hard negative mining pipelines not only accelerates iteration cycles but also ensures continuous learning from feedback on candidates sampled by the latest gradient-updating model. Furthermore, the unified multi-objective reward mechanism provides a more flexible and effective optimization strategy compared to traditional multi-loss joint training. Deployed on China’s largest e-commerce platform, Retrieval-GRPO validates its practical efficacy in improving search relevance and transaction conversion rates.

### v2

@@ 4.1.1. Dataset and Metrics
For offline evaluation, we construct three test sets: the general retrieval test set, the long-tail query evaluation set, the and long-tail experience evaluation set, all collected and filtered from Taobao’s online search logs. The model is evaluated using Hitrate@6k and Goodrate@100.
@@
Next, we provide a concise description of the offline test sets:
@@ 4.1.2. Implementation Details
The model is trained on a cluster of 256 GPUs (each equipped with 96GB memory). The training framework is implemented based on ROLL(Wang et al., 2025), an open-source distributed RL-training library optimized for large language models. During the Supervised Fine-Tuning (SFT) phase, a per-GPU batch size of 256 is configured with the AdamW optimizer ($\beta_{1}$=0.9, $\beta_{2}$=0.95), a learning rate of 1e-5, and a cosine learning rate scheduler incorporating a warmup ratio of 0.05. The architecture enforces a maximum input length of 256 tokens, with query and item embeddings projected into a 1024-dimensional L2-normalized space. MRL is applied across dimensions [1024, 2048, 3072] to enhance hierarchical representation learning. This phase completes after 1 full training epoch over the dataset.
@@ 4.1.3. Baselines
  • Query-Rewrite: Query rewriting via CSA-QR(Feng et al., 2025), followed by retrieval via inverted index.
  • Qwen-3B(Yang et al., 2025): Dense retrieval model trained based on Qwen-3B, modifying to bi-directional attention.
  • Tbstars Post-Pretrain(base): Dense retrieval model trained based on an internal closed-source Tbstars with 3B parameters, modifying to bi-directional attention.
@@ 4.2. Offline Evaluation
As shown in Table 1, the model with SFT and Retrieval-GRPO achieves optimal performance, outperforming the state-of-the-art baseline Tbstars (SFT only) with an average improvement of 4 pt across four challenging query slices. The most significant enhancement is observed in the alternative querys, showing a 4.62 pt gain. It is because the alternatives typically share identical functionality with the original items but differ in brand and price. From the lexical overlap metric perspective, queries exhibit higher overlap scores with original brands than their alternatives, bringing challenges for models relying primarily on lexical matching. For instance, a search for ”alternative PS5 controller” might incorrectly prioritize official Sony products due to brand name overlap, despite the user’s explicit request for alternatives. Through Retrieval-GRPO, wrong predictions can be corrected via relevance rewards, which helps the model overcome superficial lexical matches and develop deeper semantic understanding.
@@
Figure 2. Top-k items in R-GRPO and relevance scores assigned by the relevance model for query-item pairs
Table 6. Online Side-by-Side Human Evaluations
Query Type Case GSB Query Goodrate Item Goodrate
Q&A What medicine can make hair black? +24.30% +10.64pt +8.2pt
Alternative Alternative PS5 controller +28.57% +10.87pt +9.1pt
Overall - +29.88% +12.81pt +9.14pt
Figure 3. Distribution of Relevance Levels during R-GRPO
4.3. Ablation Study
4.3.1. Effect of R-GRPO
As shown in Table 2, we conduct ablation studies on the two-stage training framework. The SFT+Retrieval-GRPO approach significantly outperforms the SFT-only baseline. Additionally, replacing the second stage with a ranking paradigm or DPO training with offline hard negative samples both resulted in performance degradation, demonstrating the clear advantages of our proposed Retrieval-GRPO phase. Furthermore, when we removed the first-stage SFT and directly applied reinforcement learning, the performance declined substantially, indicating that while Retrieval-GRPO is effective, the first SFT stage remains essential. The training framework can be viewed as a progressive difficulty training approach: if the model lacks the basic ability to distinguish between simple positive and negative samples, it will struggle even more with challenging samples.
Figure 3 shows how the distribution of average relevance-score buckets for the top-k items evolves during training. It can be observed that L1 and L2 gradually shift towards L3 and L4 over the course of training.
Figure 4. Recall comparison between our model and the online baseline on four long-tail query-slices
4.3.2. Effect of Multi-Objective Reward
4.3.3. Effect of Reward Model
4.4. Case Study
4.4.1. Examples in R-GRPO
Through R-GRPO, the model can automatically mine hard negative samples during training and perform debiased learning based on the relevance scores. Figure 2 illustrates the top-k items selected by the dense retrieval model during R-GRPO training, along with the relevance scores assigned by the relevance model to these query-item pairs. As demonstrated in the figure, for the user search query ”square stainless steel water tank”, among the top-k items, two stainless steel square sinks get high relevance scores (0.99 and 0.99), while a plastic product (0.44) and a sink storage rack (0.29) obtain significantly lower relevance scores.
4.4.2. Case Studies on long-tail Query
Figure 2 presents a comparison between our model and the baseline currently deployed online across four types of long-tail queries. As can be seen, compared with the baseline, with R-GRPO our model can better distinguish positive samples from hard negative samples, which benefits from real-time feedback learning during the training process.
@@ 4.5. Online Evaluation
We deploy the model into the real-world pipeline and validated its effectiveness in the online search engine through A/B testing. A side-by-side evaluation was conducted between two experimental buckets, where the top 10 items for 2000 actual online queries across four challenging query slices are manually annotated using three human evaluation metrics introduced in subsection 4.1.
The specific results are shown in Table 6. We conducted fine-grained analysis on four selected challenging query slices. The most significant improvement was observed in negative-class queries, as negative semantics and literal overlap metrics typically exhibit opposing trends, posing additional challenges for models that rely on literal matching rather than deep semantic understanding for traditional dense retrieval model. Through Retrieval-GRPO, the model leverages the semantic generalization capabilities of relevance models based on LLM and employs real-time relevance feedback on retrieved top-items via reinforcement learning during training, enabling timely correction of wrong predictions.
@@ 5. Conclusion
This paper presents Retrieval-GRPO, a multi-objective reinforcement learning framework designed to address the limitations of conventional dense retrieval paradigms in e-commerce search systems. By leveraging dynamic candidate retrieval and multi-objective reward fusion in RL paradigm, our framework eliminates the dependency on labor-intensive hard negative sample mining while mitigating the seesaw effect inherent in multi-objective optimization. The proposed method integrates real-time relevance feedback from an LLM-based reward model, item quality metrics, and exclusivity constraints, and dynamically optimizes semantic representations through reinforcement learning. Extensive experiments demonstrate that Retrieval-GRPO significantly enhances semantic generalization capabilities, particularly for complex long-tail queries, by enabling real-time error correction and preference alignment during training. The framework’s elimination of static hard negative mining pipelines not only accelerates iteration cycles but also ensures continuous learning from feedback on candidates sampled by the latest gradient-updating model. Furthermore, the unified multi-objective reward mechanism provides a more flexible and effective optimization strategy compared to traditional multi-loss joint training. Deployed on China’s largest e-commerce platform, Retrieval-GRPO validates its practical efficacy in improving search relevance and transaction conversion rates.